The Oceanside Hotel is adjacent to City Coliseum, a 24,000-seat arena that is home to the city’s professional basketball and ice hockey teams and that hosts a variety of concerts, trade shows, and conventions throughout the year. The hotel has experienced the following occupancy rates for the nine years since the coliseum opened:

YEAR

OCCUPANCY RATE

1

75%

2

70 

3

72 

4

77 

5

83 

6

81 

7

86 

8

91 

9

87 

Compute an exponential smoothing forecast with α = .20, an adjusted exponential smoothing forecast with α = .20 and β = .20. Use excel and show formulas in excel.
Answer #1

Answer a: Exponential forecast, alpha= 0.2

Year Occupancy rate Exponential forecast, alpha 0.2
1 75 75.0
2 70 75.0
3 72 74.0
4 77 73.6
5 83 74.3
6 81 76.0
7 86 77.0
8 91 78.8
9 87 81.3
  • the formula to be used in Simple Exponential smoothing is

Ft+1= alpha*At + (1-alpha) Ft

At means Actual demand of t'th period, if you want to find out the Forecast through exponential smoothing= forecast of 3rd period = alpha*actual demand of 2nd period +(1-alpha) *forecast demand of 2nd period

  • remember forecast of 1st period is 75, alpha= 0.2
  • Forecast of 3rd period= alpha*A2 + (1-alpha) F2 = 0.2*70+ 0.8*75 = 74

Answer b: Trend adjusted exponential forecast, alpha= 0.2, beta= 0.2

Month, t Demand, At Exponential smoothing, Ft Trend, Tt forecast including trend, FITt
1 75 0.00 75.00
2 70 75.00 0.00 75.00
3 72 74.00 -0.20 73.80
4 77 73.44 -0.27 73.17
5 83 73.93 -0.12 73.82
6 81 75.65 0.25 75.90
7 86 76.92 0.45 77.37
8 91 79.10 0.80 79.90
9 87 82.12 1.24 83.36
  • Alpha= 0.2, Beta= 0.2
  • the formula to be used in Exponential smoothing, Ft is

Ft+1= FITt + alpha*(At+1 – FITt)

FITt is Forecast including trend, Ft+1 means Exponential forecast demand of t+1'th period, if you want to find out the Forecast of 3rd period = Forecast including trend of 2nd period + alpha*(Actual demand of 3rd period – Forecast including trend of 2nd period)

remember FIT5 = 75

  • the formula to be used in Trend smoothing, Tt is

Tt+1= Tt + beta*(Ft+1 – FITt)

Ft+1 means Exponential forecast demand of t+1'th period, if you want to find out the Trend smoothing Forecast = forecast of 3rd period = Trend smoothing Forecast of 2nd period + beta*(Exponential forecast of 3rd period – Forecast including trend of 2nd period)

remember Trend of 1st period is 0

  • the formula to be used in Forecast including Trend, FIT is

FITt= Tt + Ft